OpenStudy (anonymous):

what is the slope intercept for if to points on a (graph are 20, .91) and (40, 1.34)

OpenStudy (anonymous):

*if two

OpenStudy (anonymous):

slope intercept form is y=mx+c here slope m=(1.34-0.91)/(40-20) m=(.43/20)=0.0215 and 0.91=0.0215*20+c c=-0.48 hence the slope intercept form is y=0.0215x-0.48 i hope u understood

OpenStudy (anonymous):

kind of thanks! :)

OpenStudy (anonymous):

are u thanking me? its my pleasure,u're welcome

OpenStudy (anonymous):

yes i was thanking you! i needed it for biology and just forgot how to do it over the summer lol :p

OpenStudy (anonymous):

well done